# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ Abstract.
Lindblad dynamics and other open-system dynamics provide a promising path towards efficient Gibbs sampling on quantum computers. In these proposals, the Lindbladian is obtained via an algorithmic construction akin to designing an artificial thermostat in classical Monte Carlo or molecular dynamics methods, rather than treated as an approximation to weakly coupled system-bath unitary dynamics. Recently, Chen, Kastoryano, and Gilyén (arXiv:2311.09207) introduced the first efficiently implementable Lindbladian satisfying the Kubo–Martin–Schwinger (KMS) detailed balance condition, which ensures that the Gibbs state is a fixed point of the dynamics and is applicable to non-commuting Hamiltonians. This Gibbs sampler uses a continuously parameterized set of jump operators, and the energy resolution required for implementing each jump operator depends only logarithmically on the precision and the mixing time. In this work, we build upon the structural characterization of KMS detailed balanced Lindbladians by Fagnola and Umanità, and develop a family of efficient quantum Gibbs samplers that only use a discrete set of jump operators (the number can be as few as one). Our methodology simplifies the implementation and the analysis of Lindbladian-based quantum Gibbs samplers, and encompasses the construction of Chen, Kastoryano, and Gilyén as a special instance.
@@ 1. Introduction
For a given quantum Hamiltonian $H\in\mathbb{C}^{N\times N}$, preparing the associated Gibbs state $\sigma_{\beta}=e^{-\beta H}/\mathcal{Z}_{\beta}$ (also called quantum Gibbs sampling) has a wide range of applications in condensed matter physics, quantum chemistry, and optimization. Here $N=2^{n}$ is the dimension of the underlying Hilbert space, $\beta$ is the inverse temperature, and $\mathcal{Z}_{\beta}=\operatorname{tr}(e^{-\beta H})$ is the partition function. We assume efficient quantum access to the Hamiltonian simulation $\exp(-itH)$. The cost of a quantum algorithm is often dominated by the total Hamiltonian simulation time of $H$.
The Davies generator [Dav74, Dav76, Dav79], which is in the form of a Lindbladian (or Lindblad generator) [Lin76, GKS76], satisfies the desirable property that the Gibbs state is a fixed point of the generated dynamics, and hence can be viewed as a natural candidate for quantum Gibbs samplers. The Davies generator is typically derived as a simplified representation of weakly interacting system-bath models, following the Born-Markov-Secular [^1] The secular approximation is also referred to as the rotating wave approximation (RWA). approximation route [BP02, Lid19]. Thus its applicability range seems to be constrained by the limitations of these approximations. However, there has been a recent revival of interest in designing quantum Gibbs samplers based on the Lindblad dynamics [ML20, RWW23, CB21, CKBG23, CKG23, WT23]. These Lindbladians are constructed purely algorithmically, and may not mimic specific system-bath unitary dynamics in nature.
The key object in these Lindblad dynamics-based approaches is the following frequency-dependent jump operator
(1.1) $\widehat{A}^{a}_{f}(\omega):=\int^{\infty}_{-\infty}f(t)e^{-i\omega t}e^{iHt}A^{a}e^{-iHt}\,\mathrm{d}t,\quad\omega\in\mathbb{R}\,,$
which is a $f$-weighted Fourier transform of the Heisenberg evolution $A^{a}(t):=e^{iHt}A^{a}e^{-iHt}$ of $A^{a}$. Here $f:\mathbb{R}\to\mathbb{C}$ is a filtering function that is central to this work and will be discussed in detail, and $\{A^{a}\}_{a\in\mathcal{A}}$ is a set of (frequency-independent) coupling operators provided by the user, which represent the coupling between the system and the fictitious environment, akin to designing an artificial thermostat in classical Monte Carlo or molecular dynamics methods [FS02]. The choice of $\{A^{a}\}_{a\in\mathcal{A}}$ can be flexible and relatively simple (such as Pauli operators). The Lindblad generator for the algorithmic purpose is then formulated as
@@
(1.2) $\mathcal{L}^{{\dagger}}[\rho]=-i[G,\rho]+\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\left(\widehat{A}_{f}^{a}(\omega)\rho\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}-\frac{1}{2}\left\{\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}\widehat{A}_{f}^{a}(\omega),\rho\right\}\right)\,\mathrm{d}\omega\,.$
Here $-i[G,\cdot]$ is called the coherent part of the dynamics, and the simplest choices are $G=H$ (system Hamiltonian) or $G=0$ (no coherent term). The remaining term on the right-hand side of Eq. 1.2 is referred to as the dissipative part. In particular, the Davies generator corresponds to taking $f(t)\equiv 1$, with a carefully chosen function $\gamma:\mathbb{R}\to\mathbb{R}$, and a coherent part $-i[H,\cdot]$ (see Section 2.1).
The Gibbs state is a fixed point of the Lindblad dynamics if $\mathcal{L}^{{\dagger}}[\sigma_{\beta}]=0$. A sufficient condition to ensure this is that the generator $\mathcal{L}$ satisfies certain quantum detailed balance conditions (DBC). Here $\mathcal{L}$ and $\mathcal{L}^{{\dagger}}$ are adjoint to each other with respect to the Hilbert–Schmidt inner product. For a given precision $\epsilon>0$, we define the mixing time of the dynamics generated by $\mathcal{L}^{\dagger}$ as
@@
(1.3) $t_{\mathrm{mix}}:=\inf\left\{t\geq 0\,;\ \left\|e^{t\mathcal{L^{{\dagger}}}}\left(\rho\right)-\sigma_{\beta}\right\|_{1}\leq\epsilon,\ \forall\,\text{quantum state $\rho$}\right\}\,,$
@@
where $\left\lVert\cdot\right\rVert_{1}$ denotes the trace norm. It is worth mentioning that the quantum DBC is also instrumental for proving the finite mixing time (if it is the case), and its scaling with respect to $\beta$ and the system size [TKR+10, KT13, BCG+23, RFA24]. However, due to the non-commutativity of operators in quantum mechanics, there is no unique definition of quantum DBC [TKR+10, CM17, CM20].
@@ 1.1. Related works
Recently, [CKG23] introduced the first algorithm that requires a finite energy resolution in constructing $\widehat{A}^{a}_{f}(\omega)$, where $\mathcal{L}^{{\dagger}}$ in Eq. 1.2 exactly satisfies a less stringent version of DBC called the Kubo–Martin–Schwinger (KMS) DBC. It involves a nontrivial choice of the coherent term $G$ that is neither $H$ or 00. Under the KMS DBC, the Gibbs state $\sigma_{\beta}$ remains a fixed point of the Lindblad dynamics, and one no longer needs to keep track of the accumulated deviation from the trajectory generated by a Davies-like generator. In this setup, to prepare $\sigma_{\beta}$ to precision $\epsilon$, the integral in Eq. 1.1 can be truncated to $T=\mathcal{O}(\beta\log(t_{\mathrm{mix}}/\epsilon))$ [CKG23, Theorem I.2].
@@
The integral with respect to $\omega$ in Eq. 1.2 involves a continuously parameterized set of jump operators if $\gamma(\omega)$ is a continuous function. Although one can discretize such an integral using a quadrature scheme, the algorithm must be meticulously designed to efficiently simulate the resulting Lindblad dynamics [CW17, CKBG23, CKG23]. To the best of our knowledge, high-order Lindblad simulators designed for a finite number of jump operators, which allow for simpler implementations [LW23, DLL24], are not suitable for this task.
1.2. Contribution
In parallel to Alicki’s characterization of Lindblad generators satisfying the GNS DBC, Fagnola, and Umanità [FU07, FU10] have prescribed the necessary and sufficient conditions for a quantum Markov semigroup to satisfy the KMS DBC. This leads to a set of conditions on the jump operators and the Hamiltonian for its corresponding Lindblad generator (also see [AC21]). Building upon these works, we introduce a family of quantum Gibbs samplers satisfying the KMS DBC. This includes the construction of [CKG23] as a special instance. In particular, $\gamma(\omega)$ can be chosen to be a discrete sum of $\delta$ functions, leading to a finite number of jump operators. In fact, it is sufficient to choose $\gamma(\omega)=\delta(\omega)$, which means a single jump operator if $\left\lvert\mathcal{A}\right\rvert=1$. Our jump operators can be constructed using the standard linear combination of unitaries (LCU) routine [BCC+14, GSLW19]. As a result, our Lindblad dynamics can be efficiently simulated using any high-order simulation algorithms, including those in [LW23, DLL24]. In addition, we show that this new family of Gibbs samplers enables the selection of $f(t)$, whose Fourier transform is smooth and compactly supported. This approach simplifies the error analysis for controlling the discretization error, through the application of the Poisson summation formula. Table 1 compares the performance of a number of quantum Gibbs samplers based on the Lindblad dynamics.

### v4

@@ Abstract.
Lindblad dynamics and other open-system dynamics provide a promising path towards efficient Gibbs sampling on quantum computers. In these proposals, the Lindbladian is obtained via an algorithmic construction akin to designing an artificial thermostat in classical Monte Carlo or molecular dynamics methods, rather than treated as an approximation to weakly coupled system-bath unitary dynamics. Recently, Chen, Kastoryano, and Gilyén (arXiv:2311.09207) introduced the first efficiently implementable Lindbladian satisfying the Kubo–Martin–Schwinger (KMS) detailed balance condition, which ensures that the Gibbs state is a fixed point of the dynamics and is applicable to non-commuting Hamiltonians. This Gibbs sampler uses a continuously parameterized set of jump operators, and the energy resolution required for implementing each jump operator depends only logarithmically on the precision and the mixing time. In this work, we build upon the structural characterization of KMS detailed balanced Lindbladians by Fagnola and Umanità, and develop a family of efficient quantum Gibbs samplers using a finite set of jump operators (the number can be as few as one), akin to the classical Markov chain-based sampling algorithm. Compared to the existing works, our quantum Gibbs samplers have a comparable quantum simulation cost but with greater design flexibility and a much simpler implementation and error analysis. Moreover, it encompasses the construction of Chen, Kastoryano, and Gilyén as a special instance.
@@ 1. Introduction
For a given quantum Hamiltonian $H\in\mathbb{C}^{N\times N}$, preparing the associated Gibbs state $\sigma_{\beta}=e^{-\beta H}/\mathcal{Z}_{\beta}$ (also called quantum Gibbs sampling) has a wide range of applications in condensed matter physics, quantum chemistry, and optimization. Here $N=2^{n}$ is the dimension of the underlying Hilbert space, $\beta$ is the inverse temperature, and $\mathcal{Z}_{\beta}=\operatorname{tr}(e^{-\beta H})$ is the partition function. In this work, we assume efficient quantum access to the Hamiltonian simulation $\exp(-itH),t\in\mathbb{R}$, and we measure the cost of a quantum algorithm in terms of the total Hamiltonian simulation time of $H$.
Recently, there has been a revival of interest in designing quantum Gibbs samplers based on the Lindblad dynamics [ML20, RWW23, CB21, CKBG23, CKG23, WT23]
$\partial_{t}\rho_{t}=\mathcal{L}^{{\dagger}}\rho_{t}\,,$
which is a Markovian semigroup on quantum states. For ease of exposition, we shall refer to both the generator $\mathcal{L}^{\dagger}$ and its adjoint $\mathcal{L}$ as Lindbladian, which are related to each other via
(1.1) $\langle X,\mathcal{L}(Y)\rangle=\langle\mathcal{L}^{\dagger}(X),Y\rangle\,,\quad\text{for $X,Y\in\mathbb{C}^{N\times N}$}\,.$
[…]
@@
(1.2) $\mathcal{L}^{\dagger}_{\rm Davies}(\rho)=-i[H+H_{\rm LS},\rho]+\sum_{a,\,\omega}\gamma_{a}(\omega)\left(\widehat{A}^{a}(\omega)\rho\widehat{A}^{a}(\omega)^{\dagger}-\frac{1}{2}\left\{\widehat{A}^{a}(\omega)^{\dagger}\widehat{A}^{a}(\omega),\rho\right\}\right)\,,$
with the functions $\gamma_{a}(\omega)$ and operators $\widehat{A}^{a}(\omega)$ satisfying
$\gamma_{a}(-\omega)=e^{\beta\omega}\gamma_{a}(\omega)\,,\quad\sigma_{\beta}\widehat{A}^{a}(\omega)\sigma_{\beta}^{-1}=e^{-\beta\omega}\widehat{A}^{a}(\omega)\,,$
where $\{X,Y\}=XY+YX$ stands the anticommutator and the sum is over $a$ and $\omega$ from some discrete index sets, representing the number of couplings to the environment and Bohr frequencies, respectively. Here the commutator term involving $H+H_{\rm LS}$ is called the coherent part, and $H_{\rm LS}$ is a correction Hamiltonian called the Lamb shift satisfying $[H_{\rm LS},H]=0$. A key property for using the Davies generator as a quantum Gibbs sampler [RWW23, WT23] is that the associated dynamics $e^{t\mathcal{L}^{\dagger}_{\rm Davies}}$ admits the Gibbs state as a fixed point, equivalently, $\mathcal{L}^{\dagger}_{\rm Davies}(\sigma_{\beta})=0$, which can be directly checked from (1.2) (see Section 2.1 for a more detailed discussion).
In general, for a fixed Gibbs state $\sigma_{\beta}$, it is not an easy task to design a Lindbladian satisfying the fixed-point property $\mathcal{L}^{\dagger}(\sigma_{\beta})=0$. A sufficient condition guaranteeing this is the quantum generalization of the detailed balance condition (DBC) [KFGV77, FU07, FU10] (see 4). The quantum DBC is a set of criteria for the equilibria of quantum systems, and is crucial in understanding the dynamics and thermodynamics of such systems. Moreover, for a given precision $\epsilon>0$, we define the mixing time of a Lindblad dynamic with $\sigma_{\beta}$ as a unique fixed point by
@@
(1.3) $t_{\mathrm{mix}}:=\inf\left\{t\geq 0\,;\ \left\|e^{t\mathcal{L^{{\dagger}}}}\left(\rho\right)-\sigma_{\beta}\right\|_{1}\leq\epsilon,\ \forall\,\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0} quantum states} $\rho$}\right\}\,,$
@@
which quantifies the convergence speed of the dynamics, where $\left\lVert\cdot\right\rVert_{1}$ denotes the trace norm. The quantum DBC is also instrumental for proving the finite mixing time (if it is the case), and its scaling with respect to $\beta$ and the system size [TKR+10, KT13, BCG+23, RFA24]. However, due to the non-commutativity of operators in quantum mechanics, there is no unique definition of quantum DBC [TKR+10, CM17, CM20].
@@ 1.1. Related works
The most widely studied form of quantum DBC is in the sense of Gelfand–Naimark–Segal (GNS). The seminal result by Alicki [Ali76] characterizes the class of Lindbladians satisfying the GNS DBC, which turns out to have the same dissipative part as that of the Davies generator; see Section 2.1. The Davies generator $\mathcal{L}_{\rm Davies}$ is typically derived as a simplified representation of weakly interacting system-bath models, following the Born-Markov-Secular [^1] The secular approximation is also referred to as the rotating wave approximation (RWA). approximation route [BP02, Lid19], and its applicability range seems to be constrained by the limitations of these ideal approximations. However, for the purpose of Gibbs sampling, we may also view the Davies generator as a purely algorithmic process, without mimicking any specific system-bath unitary dynamics in nature. A much more severe issue is that the implementation of $\widehat{A}^{a}(\omega)$ requires the resolution of all Bohr frequencies, which is very challenging for most Hamiltonians (see also 7). For this reason, earlier works on the quantum thermal state preparation based on Davies generators often suffer from some unphysical assumptions, such as the rounding promise [RWW23, WT23].
In what follows, $\mathcal{A}$ denotes a discrete index set. The key object in the work [CKBG23] is the following frequency-dependent jump operator:
(1.4) $\widehat{A}^{a}_{f}(\omega):=\int^{\infty}_{-\infty}f(t)e^{-i\omega t}e^{iHt}A^{a}e^{-iHt}\,\mathrm{d}t,\quad\omega\in\mathbb{R}\,,\quad a\in\mathcal{A}\,,$
which is a $f$-weighted Fourier transform of the Heisenberg evolution $A^{a}(t):=e^{iHt}A^{a}e^{-iHt}$ of $A^{a}$. Here $f:\mathbb{R}\to\mathbb{R}$ is a real integrable filtering function, and $\{A^{a}\}_{a\in\mathcal{A}}$ is a set of (frequency-independent) coupling operators provided by the user, which represent the coupling between the system and the fictitious environment, akin to designing an artificial thermostat in the classical Monte Carlo or molecular dynamics methods [FS02]. The choice of $\{A^{a}\}_{a\in\mathcal{A}}$ can be flexible and relatively simple (such as Pauli operators). The algorithmic Lindbladian in the Schrödinger picture then takes the form
(1.5) $\mathcal{L}^{{\dagger}}(\rho)=\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\left(\widehat{A}_{f}^{a}(\omega)\rho\widehat{A}_{f}^{a}(\omega)^{\dagger}-\frac{1}{2}\Big{\{}\widehat{A}_{f}^{a}(\omega)^{\dagger}\widehat{A}_{f}^{a}(\omega),\rho\Big{\}}\right)\,\mathrm{d}\omega\,,$
@@
In the follow-up work [CKG23], Chen et al. introduced a new quantum Gibbs sampler based on (1.5), which requires a lower energy resolution in simulating $\widehat{A}^{a}_{f}(\omega)$ and hence has the reduced computational complexity, by constructing a coherent term $-i[G,\rho]$ to $\mathcal{L}^{\dagger}$ in (1.5) to make the total Lindbladian ~$\widetilde{\mathcal{L}}(X)=i[G,X]+\mathcal{L}(X)$ exactly satisfies a less stringent version of DBC called the Kubo–Martin–Schwinger (KMS) DBC. To be specific, [CKG23, Section II.C] showed that the transition part of $\mathcal{L}^{\dagger}$ in (1.5):
(1.6) $\mathcal{T}(\rho)=\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\widehat{A}_{f}^{a}(\omega)\rho\widehat{A}_{f}^{a}(\omega)^{\dagger}\,\mathrm{d}\omega$
satisfies the $\sigma$-KMS DBC if $f$ and $\gamma$ are Gaussians with some parameter constraints. Furthermore, if $f(t)$ is chosen to be a real function, then the Gaussian function turns out to be the unique choice (up to a linear combination of $\gamma$, see [CKG23, Appendix D]). On the other hand, [CKG23, Lemma II.1] showed one can always design a coherent term $G$ such that the superoperator $\mathcal{R}(\rho)-i[G,\rho]$ is KMS detailed balanced, where $\mathcal{R}(\cdot)$ is the decay part in (1.5):
(1.7) $\mathcal{R}(\rho)=-\frac{1}{2}\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\Big{\{}\widehat{A}_{f}^{a}(\omega)^{\dagger}\widehat{A}_{f}^{a}(\omega),\rho\Big{\}}\,\mathrm{d}\omega\,.$
Thanks to the exact KMS DBC, the Gibbs state $\sigma_{\beta}$ remains a fixed point of the Lindblad dynamics, and one no longer needs to keep track of the accumulated deviation and distinguish the energy levels of $H$ to very high precision. In this setup, it was proved [CKG23, Theorem I.2] that to prepare $\sigma_{\beta}$ to precision $\epsilon$, the integral in Eq. 1.4 can be truncated to $T=\mathcal{O}(\beta\log(t_{\mathrm{mix}}/\epsilon))$ and the cost can be significantly reduced.
@@
It is worth pointing out that for each jump proposal $A^{a}$, the integral with respect to $\omega$ in Eq. 1.5 involves a continuously parameterized set of jump operators if $\gamma(\omega)$ is a continuous function. Although one can discretize such an integral using a quadrature scheme, the algorithm must be meticulously designed to efficiently simulate the resulting Lindblad dynamics [CW17, CKBG23, CKG23]. To the best of our knowledge, high-order Lindblad simulators designed for a finite number of jump operators, which allow for simpler implementations [LW23, DLL24], are not suitable for this task.
Along with [CKBG23, CKG23], there are other algorithms [RWW22] for the thermal state preparation also aiming to approximate the GNS detailed balanced Lindbladian (i.e., the Davies generator). Similar to [CKBG23], the truncation time of these methods often depends polynomially on $1/\epsilon$, resulting in the total cost scales polynomially with $1/\epsilon$. Table 1 summarizes the total costs of the algorithms in [CKBG23, CKG23] and some earlier related works. A more comprehensive overview of quantum Gibbs samplers can be found in [CKBG23] and references therein.
In parallel to Alicki’s characterization of Lindblad generators satisfying the GNS DBC, Fagnola, and Umanità [FU07, Theorems 7.2 and 7.3] have prescribed the necessary and sufficient conditions for a Lindblad dynamics to satisfy the KMS DBC. This leads to a set of conditions on the jump operators and the Hamiltonian for its corresponding Lindblad generator; see also [FU10, AC21]. However, it is unclear from the results in [FU07] how to choose suitable jump operators for the quantum thermal state preparation, and they also did not provide an explicit form of the coherent term required for the KMS detailed balance. In this sense, [CKG23, Lemma II.1] shares some commonalities to the results in [FU07] and provides a more explicit form of the Lindbladian.
1.2. Contribution and outline
In this work, we first provide a detailed introduction to the canonical forms of Lindbladians with various detailed balance conditions in Section 2. In particular, in Section 2.2, we revisit the arguments in [FU07], along the lines of the recent work [AC21], and provide a self-contained construction for the $\sigma_{\beta}$-KMS detailed balanced Lindbladian with explicit formulations of the jumps and the coherent term (see 10 and Eq. 3.1). This is the starting point of our quantum Gibbs samplers constructed and analyzed in Section 3.
The idea underlying our quantum Gibbs sampler is analogous to the classical Markov chain-based sampling. In the classical context, a state undergoes a candidate jump according to a transition matrix. The proposed new state is then accepted with a certain probability to ensure that the resulting modified transition matrix satisfies a detailed balance. A natural question that arises when designing efficient quantum Gibbs samplers is:
How to modify a predefined set of jump proposals (i.e., coupling operators) such that the resulting Lindblad dynamics satisfies the quantum detailed balance for the target Gibbs state and meanwhile can be efficiently simulated on a quantum computer?
Based on the structural characterization in 10, we provide an affirmative answer to this question. To the best of our knowledge, our approach is the first framework that enables the use of general filter functions to modify a set of predefined coupling operators, resulting in dynamics that can be efficiently simulated and exactly satisfy the KMS detailed balance condition.
In contrast to [CKBG23, CKG23], which can be regarded as smoothened approximations of Davies generators, our new framework provides significantly more flexibility in designing quantum Gibbs samplers. Recall from [CKG23, Appendix D] that to make the transition part (1.6) satisfy the KMS DBC, the filtering function $f$ in (1.4) has to be a Gaussian if it is assumed to be real, and then the outer integration in (1.6) with respect to a continuous function $\gamma(\omega)$ is necessary. Intuitively, this is because when $f$ is real, its Fourier transform must be conjugate symmetric. We need an additional integration to induce the biased energy transition. However, if the main goal is to ensure the KMS detailed balance (instead of approximating Davies generators), by considering a general class of filtering functions with a mild symmetry condition (see (1.8)–(1.9) below), we find that $f$ can be relaxed to be complex and the outer integration in $\omega$ can be removed.
To be specific, in Section 3.1, we show that the admissible jump $L$ for KMS detailed balance is of the form: for a self-adjoint operator $A$ with $A_{\nu}=\sum_{\lambda_{i}-\lambda_{j}=\nu}P_{i}AP_{j}$,
$L=\sum_{\nu\in B_{H}}e^{-\beta\nu/4}A_{\nu}\,,$
where $B_{H}$ is the set of Bohr frequencies (see (2.11)–(2.13) for related definitions). We may overparameterize it by adding a weighting function $q(\nu)$:
(1.8) $L=\sum_{\nu\in B_{H}}\widehat{f}(\nu)A_{\nu}\,,\quad\widehat{f}(\nu)=e^{-\beta\nu/4}q(\nu)\,,$
with $q$ being conjugate symmetric $q(-\nu)=\overline{q(\nu)}$, and then, without loss of generality, we can assume the operator norm $\|A\|\leq 1$. Here, intuitively, the Fourier transform of the filtering function $\widehat{f}$ helps modify the transition probability between eigenstates with an energy difference $\nu$ to satisfy the KMS detailed balance. This jump operator can be efficiently simulated in the time domain:
(1.9) $L=\int_{-\infty}^{\infty}f(t)e^{iHt}Ae^{-iHt}\,\mathrm{d}t\quad\text{with}\quad f(t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\widehat{f}(\nu)e^{-i\nu t}\,\mathrm{d}\nu\,,$
when $f(t)$ decays rapidly as $\left\lvert t\right\rvert\to\infty$ (see also (2.15)–(2.16)). This is the main motivation for us to consider the Gevrey function class with compact support in 13 as weighting functions $q(\nu)$ [AHR17, HR19], whose Fourier transform has the desired sub-exponential decay as $|t|\to\infty$ (see Lemmas 29 and 30), so that one could approximate the integral (1.9) accurately with a moderate truncation time $\widetilde{\mathcal{O}}(\beta\mathrm{log}^{1+o(1)}(t_{\rm mix}/\epsilon))$ (see 15 and 18). Once we design a family of efficient jumps $\{L_{a}\}$ via $\{(A_{a},f^{a})\}$, the coherent term can be readily defined via Eq. 2.33 in 10, admitting a similar time integral formulation as (1.9) (see Eqs. 3.10 and 3.11). The quantum algorithm for the proposed quantum Gibbs sampler is discussed in detail in Section 3.3. In Section 4, we elaborate on how to fit the construction of [CKG23] into our algorithmic framework as a special instance.
In addition, we emphasize that compared to the algorithm in [CKG23], our Gibbs sampler only uses one jump operator per jump proposal $A^{a}$. In particular, we can directly achieve a Metropolis-type transition using our framework while exactly satisfying the KMS DBC (see Fig. 1). This greatly simplifies the quantum simulation process. Thanks to 15 built on the technical 13 via the Poisson summation formula, our jump operators can be constructed using the standard linear combination of unitaries (LCU) routine [BCC+14, GSLW19]. As a result, our Lindblad dynamics can be efficiently simulated using any high-order simulation algorithms, including those in [LW23, DLL24]. We show in 18 that our approach and that in [CKG23] have a comparable computational cost when the best available Lindblad simulation algorithms with near-optimal costs are used. One should note that 13 (Gevrey-type weighting functions) is mainly for simplifying the discretization error analysis and not essential for the design of our quantum Gibbs sampler. We believe that a broader class of smooth functions can also serve as weighting functions without affecting the total quantum algorithm cost.